The emf of the driver cell of a potentiometer is 2V & its internal resistance is negligible . The length of the potentiometer wire is 100cm & resistance is 5 ohm.How much resistance is to be connected in series with the potentiometer wire to have a potential gradient of 0.05V/cm?

Grade:11

5 Answers

Arjun Shamsunder
11 years ago

Potential gradient K= 0.05 V/cm = 5V/m
Driver cell emf E=2V
Resistance of potentiometer wire R1= 5Ω/m
Series resistance, Rs= ?

Now, to have K=5V/m, current flowing through the wire must be 1 Ampere (  since current I= 5V/5Ω=1A)
So, for the driver cell of emf 2V to generate 1A of current, series resistance ro be connected is 2Ω( since Rs= E/I=2/1=2Ω)

Therefore, Rs=2Ω

Anubhav Sharma
11 years ago

the resistance connected to be in series wih wire should be 7.5 ohm

is it right 

Anubhav Sharma
11 years ago

the resistance to be connected should be 7.5 ohm 

is it right

Ranjita yadav
11 years ago

The ans is 1995ohm.

Arjun Shamsunder
11 years ago

I think its a wrong question